OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? a. x − 3y = 9 b. x + 3y = 9 c. 3x − y = −9 d. 3x + y = 9

OpenStudy (saifoo.khan):

Step1 = Find slope. Step2 = Use the formula, y-y1=m(x-x1)

OpenStudy (anonymous):

b.x+3y=9

OpenStudy (anonymous):

\[y\text{ =}-\frac{x}{3}+3\]

OpenStudy (precal):

 step 1